OpenStudy (anonymous):

add 6/w + 1/w^2=7/w^2??? add 6/w + 1/w^2=7/w^2??? @Mathematics

OpenStudy (anonymous):

w=-1/6

OpenStudy (anonymous):

no wait is it 1/6

OpenStudy (anonymous):

you can't add them, denominators are not the same

OpenStudy (anonymous):

yea i keep wanting to add them

OpenStudy (anonymous):

no wait i have to have a w dont I?

OpenStudy (turingtest):

\[{6 \over w}+{1 \over w^2}={7 \over w^2}\] multiply first term by w/w: \[{6w \over w^2}+{1 \over w^2}={7 \over w^2}\] multiply all by w^2 to get rid of the denominator:

OpenStudy (anonymous):

so was the correct answer what i had?

OpenStudy (turingtest):

6w+1=7 6w=6 w=1

OpenStudy (turingtest):

If that was the answer you had...

OpenStudy (anonymous):

yea i came up with 7/w^2

OpenStudy (turingtest):

Well that's what the right side turns into...

OpenStudy (turingtest):

Oh wait, it was always that.

OpenStudy (turingtest):

do you understand what I did multiplying the first expression by w/w?

OpenStudy (anonymous):

yea i see what your saying. thanks.